The width of a rectangle is 6 ft less than the length. The area of the rectangle is 112 ft^2. Find the length L and width W of the rectangle.
Answer by ikleyn(52776)

.

Let "w" be the width, in feet.

Then the length is  (w+6)  ft.


The area equation is


    w*(w+6) = 112.


Simplify and find w

    
    w^2 + 6w - 112 = 0


    (w+14)*(w-8) = 0.


The roots are -14 and 8.

You need a positive root, which is 8.


ANSWER.  The width is  8 ft;  the length is  8+6 = 14 ft.

Solved.